Tutorials:

Mean, Range, Median, and Mode Tutorial Creator of Tutorial: AlexandervonHumboldt2

Tutorials:

Mean, Range, Median and Mode Tutorial! \(\Huge\color{red}{ ☞ }\) What is mean? Mean is the average of the set of numbers. To find the mean add all the values together and divide by the number of values in the set. \[\frac{ x_1+x_2+x_3+x_4+...+x_n }{ n }=mean\] Examples: Lets find mean of the following sets: {1, 5, -8, -65, 85, 12, 32} \[\frac{ 1+5+(-8)+(-65)+85+12+32 }{ 7 }=62/7\] Another example: {-1, 2, 10, 25, -30} \[\frac{ -1+2+10+25-30 }{ 5 }=6/5\] \(\Huge\color{red}{ ☞ }\) What is range? Range is the diference between the biggest number and the smallets number in the set. To find range we need to subtract smallest number from biggest number. Let x_1 be the smallest and x_6 be the biggest. \[{x_1, x_2, x_3, x_4, x_5, x_6}\] The range is \[x_6-x_1\] Example: {8, -9, 10, 2, 3, 65, 85} The biggest number is 85. The smallest number is 2. Thus we subtract 2 from 85: 85-2=83 83 is the range. \(\Huge\color{red}{ ☞ }\) What is Mode? Mode is the number that has the most frequency in the set. To find mode you simplify need to find which number has the most frequency. Example: {1, 2, 3, 98, 66, 85, 99, 99, 45, 99, 32} The most frequency has 99. Thus 99 is the mode. \(\Huge\color{red}{ ☞ }\) What is median? Median is the middle of the set of numbers. To find it you need to order the number from least to greatest and find the misddle of the set. Example: {1, 2, -8, -6, 12, 17, 8} Order them: -8, -6, 1, 2, 8, 12, 17 The middle is 2. Thus median is 2. NOTE! These can be a set which has even number of values. Like {1, 2, 3, 4, 5, 6} In such case you need to find the average of the 2 medains: The meddians are 3 and 4. Now find the average: (3+4)/2=7/2=3.5 Thus 3.5 is the median. Now it is your turn! \(\Huge\color{red}{ ☞ }\) Find the mean, median, mode and range of the following set: {1, -8, -8, 99, 67, 58, 10, -23, -23, 10}

Ashely:

THANKS